OpenStudy (anonymous):

Can someone help me out?? I need the y-intercept and the x-intercept of the equation: 4x-5y=-6

OpenStudy (anonymous):

finding x, set y to 0 and solve: 4x-5(0)=-6 4x=-6 x=-3/2 vice versa 4(0)-5y=-6 -5y=-6 y=-6/-5

OpenStudy (anonymous):

Thanks alot!